a boat travels 2 km upstream in 2 km back. The time for the round trip is 7 hours.The speed of the stream is 2 km/hr. What is the speed of the boat in still water

Answer by mananth(16946)

let speed of boat in still water be x kmph
speed of current = 2 kmph
..
upstream speed = x-2
downstream speed = x+2
..
t= d/r
time upstream + time downstream = 7 hours
distance = 2km
2/(x-2)+2/(x+2)=7
2(x+2)+2(x-2)/(x+2)(x-2)=7
2x+4+2x-4 = 7(x+2)(x-2)
4x=7(x^2-4)
4x=7x^2-28
7x^2-4x-28 =0
solve using quadratic equation
discriminant = 800
x1= (+4+sqrt(800))/14
x1= 2.3 km/h speed of boat in still water
x2 is negative so ignore
...
m.ananth@hotmail.ca